OpenStudy (anonymous):

Find the solutions of each system of equations. 1. x^2 + y^2 =16 x^2 / 16 - y^2/ 9 =1

OpenStudy (anonymous):

x^2 + y^2 =16.................i x^2 / 16 - y^2/ 9 =1.............ii

OpenStudy (anonymous):

Now, from i : x^2=16-y^2

OpenStudy (anonymous):

substitute that in eqn ii and solve for y

OpenStudy (anonymous):

can u?